OpenStudy (anonymous):

Can somebody please help? I need help on this one problem 6x – 4y = 54 –9x + 2y = –69 please show work

OpenStudy (anonymous):

you might notice if you multiply the second eq and ten add it to the first, something magical happens. :)

OpenStudy (anonymous):

I'm not sure what you are trying to do to the equations, but if you are trying to find x and y, here's how: 6x-4y=54 add 4y to both sides 6x=54+4y divide by 6 on both sides x=9+2/3y plug x in to the other equation: -9(9+(2/3)y)+2y=-69 distribute the -9 -81-6y+2y=-69 combine like terms -81-4y=-69 add 81 to both sides -4y=12 divide -4 on both sides y=-3 plug y into the other equation: 6x-4(-3)=54 multiply 6x+12=54 subtract 12 from both sides 6x=42 divide by 6 on both sides x=7